1. A fund had $60 million under management at the beginning of January. The fund's return for January was -1.9% and the fund received $4.3 million new capital inflow at the month end. What was the fund's portfolio value at this point?


2. Continued with Q1. Suppose the fund earned 5.6% return but experienced $5.3 million capital outflow for February. What was its portfolio value at the end of Feb. ?


3. Continued with Q1 and Q2. For March the fund earned 4.8% and there was no new inflow. What was the fund's portfolio value at the end of March?


4. Continued with Q1-Q3. What is the fund's money weighted average return?

5. Continued with Q1-Q5. What is the geometric average monthly return of the fund for the three months?
Answer #1

1) Jan Value = 60 x (1 - 1.9%) + 4.3 = $63.16 million

2) Feb Value = 63.16 x (1 + 5.6%) - 5.3 = $61.40 million

3) Mar Value = 61.40 x (1 + 4.8%) = $64.34 million

4) Money weighted average return can be calculated using IRR function on a calculator

Insert CF0 = -60, CF1 = 4.3, CF2 = -5.3, CF3 = 64.34

=> Compute IRR = 1.86% is the money weighted average return.

5) Geometric average monthly rate = ((1 - 1.9%) x (1 + 5.6%) x (1 + 4.8%))^(1/3) - 1 = 2.78%